dielectric with a positive charge of 5 C has 18.75 × electrons added to it. What is the net charge

1
Expert's answer
2022-01-25T03:04:01-0500
  • charge on electrons=- 1.60217662 × 10-19 coulombs
  • charge on 18.75 electrons=- 1.60217662 × 10-19*18.75= - 30.04 *10-19coulombs
  • Net charge= (5 - 30.04 *10-19) coulombs